OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 1/2 x - 2y = 4 and 4y - x = -8

OpenStudy (anonymous):

First multiply by 2 to get rid of the fraction. You'll get: x-4y = 8 Make y positive by multiplying by -1 4y-x=-8 Second Equation : 4y-x=-8 Both are the same equations therefore the same line. Solution should be all real numbers.